# Sylva: Tailoring Personalized Adversarial Defense in Pre-trained Models via Collaborative Fine-tuning

### v1

@@ 2.4.2. Impact of the Personalized Framework
Figure 5. Comparison of aggregation methods for adversarial training in heterogeneous environments
2.4.3. Impact of the Aggregation Method
@@
Many studies have shown that data heterogeneity in federated adversarial training causes model drift on client devices, with direct aggregation slowing convergence (Zhou et al., 2022; Chen et al., 2022; Li et al., 2023a). We believe that, in the personalized federated adversarial framework, LoRA parameters may also be impacted by non-IID data, leading to drifts that undermine the adversarial generalization of the fine-tuned backbone.
@@
We simulate a system with 15 clients, where each client’s heterogeneous data follows a Dirichlet distribution (Li et al., 2020a). After 10 epochs of local adversarial training, we apply PCA algorithm to the LoRA parameters, reducing them to 3 dimensions, as shown in Fig. 3. The results reveal drifts in some local LoRA models due to data heterogeneity. We design two aggregation strategies: one aggregates all models (All), while the other selectively aggregates only those whose PCA-compressed parameters are close, excluding models with significant drifts (Sel). After the same federated adversarial fine-tuning, the averaged test results for each client, shown in Fig. 4, demonstrate that only aggregating normal models yields higher AR and BA metrics. This strategy removes models with excessive drift, thereby accelerating the generalization and convergence of the adversarial backbone.
Discovery 3: Data heterogeneity can cause LoRA model drift, and models with excessive drift, when aggregated, can enhance convergence speed and reduce adversarial training performance.
However, the drifted model may contain unique knowledge not present in others. Discarding it outright could expedite convergence but risks losing valuable information, ultimately undermining the effectiveness of adversarial training. Moreover, compressing all LoRA parameters using PCA is computationally intensive, with efficiency significantly declining as the number of clients scales. Thus, a more refined aggregation approach is needed.
2.4.4. Limitations of the Robustness-Accuracy Trade-off
@@
In adversarial training within pre-trained models, the key challenge is to maximize benign accuracy while maintaining adversarial robustness. The method in (Zhou et al., 2024) enhances benign accuracy by first computing the loss for each model layer using a set of adversarial samples. The top-$p$ layers with the lowest loss—indicating minimal sensitivity to robustness—are selected for unfreezing, while the others remain frozen. The model is then trained on benign data, enhancing benign accuracy without sacrificing robustness.
@@
However, the method has a critical flaw: the $p$ selected layers may not remain optimal after joint selection due to layer interactions affecting overall performance. To verify this, we conduct the following experiment: for the robust models trained on the 15 clients, we measure the loss of each layer as in (Zhou et al., 2024). We then select multiple combinations of $p$ layers, unfreezing them while freezing the others, and retrained the model using benign data.
@@
Figure 6. The overview of Sylva. (1) Left: Phase 1—Clients and the cloud collaborate in personalized federated adversarial fine-tuning to obtain a generalized backbone and personalized classifier. (2) Right: Phase 2-Clients apply a game-based algorithm to fine-tune layers, enhancing robustness while improving accuracy.
3. Methodology
@@
In this section, we present the framework of the Sylva algorithm, offering a comprehensive explanation of its key components and the overall training process.
3.1. Overview
@@
Fig. 6 provides an overview of the framework and training process of the Sylva algorithm, which is divided into two distinct phases.
@@
In the Phase 1, personalized adversarial fine-tuning, the algorithm is implemented within a federated learning framework, where each client employs the LoRA fine-tuning method to personalize adversarial training. Each client fine-tunes its pre-trained model using our proposed novel loss function, updating only the LoRA and classifier modules. After local training rounds, LoRA parameters are uploaded to the cloud, where a ball-tree-based method assigns weights for aggregation and updates. Clients then download the aggregated LoRA model, iterating this process until convergence. Through this method, each client ultimately obtains a shared global backbone, which incorporates the aggregated LoRA modules. This backbone is capable of generalizing robust feature extraction from adversarial samples. Simultaneously, each client retains its personalized classifier module, which ensures targeted defense against adversarial attacks specific to its own data distribution.
@@
The Phase 2, cooperative game-based fine-tuning, aims to enhance the benign accuracy of models on standard downstream tasks while preserving their robustness. For the personalized models trained by each client in Phase 1, we assume they already exhibit sufficient robustness against adversarial samples. However, their inference performance on local benign data can still be further improved. To achieve this, each client independently and in parallel fine-tunes its model using its local benign dataset. In this phase, we introduce a value function and adopt a cooperative game-based approach leveraging Shapley values to identify the optimal layers of the large model to freeze during training. This strategic layer selection maximizes the value function, allowing the models to improve their accuracy on benign data while incurring minimal loss of robustness. The outcome is a refined client-specific model that balances robust feature extraction with enhanced performance on standard tasks.
3.2. Personalized Adversarial Fine-tuning
@@
We introduce the training details of Phase 1, including the design of the novel LoRA-based training paradigm, the optimization of the aggregation algorithm, and the workflow of training process.
3.2.1. Adaptive Class-balanced Dynamic Weighted Loss
@@
Current adversarial training algorithms predominantly focus on designing loss functions that balance the relationship between benign data and adversarial samples, seeking an optimal trade-off between accuracy and robustness. Building on this, we aim to design a novel loss function specifically adapted to LoRA fine-tuning, incorporating a class-weighting strategy that accounts for heterogeneous data distributions, enhancing adversarial training for each client’s unique local data.
@@
(5) $h_{i_{c}}^{Base}=\frac{1-\gamma}{1-\gamma^{n_{i_{c}}}},$
@@
(6) $h_{i_{c}}^{Ada}(t)=\epsilon\cdot h_{i_{c}}^{Ada}(t-1)+(1-\epsilon)\cdot h_{i_{c}}^{Base}(t),$
@@
(8) $\mathcal{L}_{A}(w_{i},x^{adv},y)=\min\sum_{(x,y)\in\mathcal{D}_{i}}h_{i_{y}}\cdot\mathcal{L}_{CE}(\mathcal{F}(w_{i},x^{adv}),y),$
@@
(9) $\mathcal{L}_{S}(w_{i},x,x^{adv})=\mathcal{L}_{KL}(\mathcal{F}(w_{i},x),\mathcal{F}(w_{i},x^{adv}))$
@@
For the backbone module of pre-trained models, our optimization objective is to improve its generalization in feature extraction for adversarial samples. Meanwhile, the classifier module focuses on personalized classification performance. To account for client heterogeneity, we introduce a regularization loss to prevent excessive LoRA updates and preserve backbone generalization. The loss function is formulated as
(10) $\mathcal{L}_{R}(w_{i}^{L},w_{g}^{L})=\left\|w_{i}^{L}-w_{g}^{L}\right\|_{2}^{2},$
@@
where $w_{i}^{L}$ and $w_{g}^{L}$ represent the LoRA parameters of client $i$ and the global aggregated LoRA parameters, respectively. Finally, by combining Eq. 8, Eq. 9 and Eq. 10, we weight the above loss terms to obtain the adversarial training loss as
(11) $\mathcal{L}=\mathcal{L}_{A}+\lambda_{1}\mathcal{L}_{S}+\lambda_{2}\mathcal{L}_{R}.$
@@
Each client conducts local adversarial training based on Eq. 11, with updates restricted to the LoRA and classifier modules. Clients upload LoRA parameters after fixed epochs for global aggregation on the server.
3.2.2. Ball-tree-based Aggregation
Upon receiving the LoRA parameters from each client, the server faces model heterogeneity due to variations in local data distributions. Since the primary goal of adversarial training for the pre-trained model’s backbone, fine-tuned with LoRA, is to generalize feature extraction from adversarial samples, models trained on highly biased client data could potentially disrupt the global model. Uniform aggregation of model weights across all clients may, therefore, hinder convergence efficiency.
We adopt a ball-tree-based (Dolatshah et al., 2015) retrieval method to assign aggregation weights to each client model. The LoRA model parameters received from each client are flattened into vectors, represented as
(12) $\mathbf{G}=\{\mathcal{G}(w_{1}^{L}),\mathcal{G}(w_{2}^{L}),\ldots,\mathcal{G}(w_{N}^{L})\},$
(13) $d_{ij_{m}}=\|\mathcal{G}(w_{i}^{L})-\mathcal{G}(w_{j_{m}}^{L})\|_{2},\quad m=1,2,\ldots,k$
(14) $q_{i}=\frac{\sum_{m=1}^{k}\exp\left(-\frac{d_{ij_{m}}}{\sigma^{2}}\right)}{\sum_{i=1}^{N}\sum_{m=1}^{k}\exp\left(-\frac{d_{ij_{m}}}{\sigma^{2}}\right)}.$
The weighting scheme favors clients with stable data, centrally located in high-dimensional space and exhibiting stronger robust generalization, while assigning lower weights to those with more extreme, dispersed distributions. Finally, by combining Eq. 3, the server can obtain the aggregated LoRA model as
(15) $w_{g}^{L}=\sum_{i=1}^{N}\frac{q_{i}|D_{i}|w_{i}}{\sum_{i=1}^{N}q_{i}|D_{i}|}.$
Our aggregation computes weights over high-dimensional parameters, where naive nearest neighbor search incurs $O(N^{2})$ complexity. Leveraging the ball tree algorithm reduces this to $O(N\log N)$ by efficiently capturing distance relations, enabling scalability to large-scale scenarios.
3.3. Cooperative Game-based Fine-tuning

### v3

@@ II-D2 Impact of the Personalized Framework
Figure 5: Comparison of aggregation methods for adversarial training in heterogeneous environments
II-D3 Impact of the Aggregation Method
@@
Many studies show that data heterogeneity in federated adversarial training causes model drift on clients, with direct aggregation slowing convergence [46, 4]. In the personalized federated adversarial framework, LoRA-FA parameters may likewise be affected by non-IID data, leading to drifts that weaken the adversarial generalization of the fine-tuned backbone.
@@
We simulate a system with 15 clients whose heterogeneous data follow a Dirichlet distribution, and include a Byzantine client performing label-flipping during training. After 10 epochs of local adversarial training, we apply PCA algorithm to the LoRA-FA parameters, reducing them to 3 dimensions, as shown in fig. 3. The results reveal drifts in some local LoRA-FA models due to data heterogeneity, while the Byzantine client also drifts and overlaps with some honest heterogeneous models. We design two aggregation strategies: one aggregates all models (All), while the other selectively aggregates only those whose PCA-compressed parameters are close, excluding models with significant drifts and Byzantine attacks (Sel). In this setting, naive drift-based aggregation may exclude honest heterogeneous models. After the same federated adversarial fine-tuning, the averaged test results for each client, shown in fig. 4, demonstrate that only aggregating normal models yields higher AR and BA metrics. This motivates a Byzantine-aware aggregation strategy that preserves heterogeneity-aware adversarial training.
Discovery 3: Under data heterogeneity, both honest and Byzantine clients can exhibit drift, making naive aggregation insufficient for personalized adversarial training.
However, drifted models may still encode client-specific knowledge arising from data heterogeneity, while similar drift patterns can also be induced by Byzantine updates, and indiscriminately discarding such models can accelerate convergence at the cost of losing valuable information for adversarial training.
II-D4 Limitations of the Robustness-Accuracy Trade-off
@@
In adversarial training within pre-trained models, the key challenge is to maximize benign accuracy while maintaining adversarial robustness. The method in [48] enhances benign accuracy by first computing the loss for each model layer using a set of adversarial samples. The top-$p$ layers with the lowest loss—indicating minimal sensitivity to robustness—are selected for unfreezing, while the others remain frozen. The model is then trained on benign data, enhancing benign accuracy without sacrificing robustness.
@@
However, the method has a critical flaw: the $p$ selected layers may not remain optimal after joint selection due to layer interactions affecting overall performance. To verify this, we conduct the following experiment: for the robust models trained on the 15 clients, we measure the loss of each layer as in [48]. We then select multiple combinations of $p$ layers, unfreezing them while freezing the others, and retrained the model using benign data.
@@
Figure 6: Overview of Lorica. (1) Phase 1—Clients and the server collaborate in federated adversarial fine-tuning to obtain a generalized backbone and personalized classifier. (2) Phase 2—Clients fine-tune layers using a forward-gating algorithm, improving benign accuracy while preserving robustness.
III Methodology
@@
In this section, we present the framework of the Lorica algorithm, offering a comprehensive explanation of its key components and the overall training process.
III-A Overview
@@
fig. 6 provides an overview of the framework and training process of the Lorica algorithm, which is divided into two distinct phases.
@@
In the Phase 1, personalized adversarial fine-tuning, the algorithm is implemented within a federated learning framework, where each client employs the LoRA-FA fine-tuning method to personalize adversarial training. Each client uploads its LoRA-FA parameters to the server, where a Byzantine-aware ball-tree-based aggregation method generates both a global model and expert models. Clients then download the respective models and perform personalized local fine-tuning using an MoE-inspired adaptive optimization strategy. This ensures robust global generalization while maintaining client-specific personalization. Through this method, each client obtains a shared global backbone integrating the aggregated LoRA-FA modules, enabling robust feature extraction from adversarial samples. Simultaneously, clients retain personalized classifier modules to ensure targeted defense against adversarial attacks specific to their data distributions.
@@
The Phase 2, cooperative game-based fine-tuning, aims to enhance the benign accuracy of models on standard downstream tasks while preserving their robustness. For the personalized models trained by each client in Phase 1, we assume they already exhibit sufficient robustness against adversarial samples. However, their inference performance on local benign data can still be further improved. To achieve this, each client independently and in parallel fine-tunes its model using its local benign dataset. In this phase, we propose a forward-gating network that adaptively selects which model layers to freeze or fine-tune during training. Guided by an optimization objective designed to balance benign accuracy and adversarial robustness, the network fine-tunes layers with the highest improvement potential, enhancing accuracy on benign data while minimizing robustness loss.
III-B Personalized Adversarial Fine-tuning
@@
We introduce the training details of Phase 1, including the optimization of the aggregation algorithm and the design of the novel LoRA-based training paradigm.
III-B1 Ball-tree-based Aggregation
Similar to other distributed algorithms, after clients upload their models, the server is responsible for aggregating the received parameters. In this case, the server aggregates the uploaded LoRA-FA parameters to construct a global model. This aggregation process is critical for ensuring the generalization of the model while addressing the challenges posed by heterogeneous client data distributions.
We adopt a ball-tree-based [8] retrieval method to assign aggregation weights to each client model. The LoRA-FA model parameters received from each client are flattened into vectors, represented as
$\mathbf{G}=\{\mathcal{G}(w_{1}^{L}),\mathcal{G}(w_{2}^{L}),\ldots,\mathcal{G}(w_{N}^{L})\},$ (5)
$d_{ij_{m}}=\|\mathcal{G}(w_{i}^{L})-\mathcal{G}(w_{j_{m}}^{L})\|_{2},\quad m=1,2,\ldots,k$ (6)
$q_{i}=\frac{\sum_{m=1}^{k}\exp\left(-\frac{d_{ij_{m}}}{\sigma^{2}}\right)}{\sum_{i=1}^{N}\sum_{m=1}^{k}\exp\left(-\frac{d_{ij_{m}}}{\sigma^{2}}\right)}.$ (7)
The weighting scheme favors clients with stable data, centrally located in high-dimensional space and exhibiting stronger robust generalization, while assigning lower weights to those with more extreme, dispersed distributions.
Our aggregation computes weights over high-dimensional parameters. Naive nearest neighbor search has $O(N^{2})$ complexity, but the ball tree algorithm reduces it to $O(N\log N)$, enabling efficient and scalable distance computation. Leveraging the ball tree’s inherent hierarchical clustering and fast nearest-neighbor retrieval, we partition the tree at depth $d$ into $E$ clusters, each maintaining an expert model, where $e(i)$ denotes the index of the cluster to which client $i$ belongs. To identify Byzantine clients under non-IID data distributions, we aim to detect the most distributionally deviant clients within each cluster. Let the set of clients in the $e$-th cluster be denoted by $\mathbf{C}_{e}$. We define the robust cluster center as $\mu_{e}=\mathrm{median}\{\mathcal{G}(w_{i}^{L})\mid i\in\mathbf{C}_{e}\}$, where the median is computed element-wise. The distance of each client update to the cluster center is then given by $\psi_{i}=\|\mathcal{G}(w_{i}^{L})-\mu_{e}\|_{2}$. During aggregation, client updates can be filtered and reweighted according to these distances as
$\tilde{q}_{i}=q_{i}\cdot\mathbf{1}\{\psi_{i}\leq\tilde{\psi}_{e(i)}+\kappa\operatorname{MAD}_{e(i)}\},$ (8)
where $\tilde{\psi}_{e}=\mathrm{median}\{\psi_{i}\mid i\in\mathbf{C}_{e}\}$, $\mathrm{MAD}_{e}=\mathrm{median}\{|\psi_{i}-\tilde{\psi_{e}}|\mid i\in\mathbf{C}_{e}\}$. Finally, by combining eq. 3, the server can obtain the aggregated LoRA-FA model as
$w_{g}^{L}=\sum_{i=1}^{N}\frac{\tilde{q_{i}}|D_{i}|w_{i}^{L}}{\sum_{i=1}^{N}\tilde{q_{i}}|D_{i}|}.$ (9)
Within each cluster, the LoRA-FA parameters are averaged to produce the expert $w_{e}^{L}$ by a robust trimmed mean strategy. The server transmits both the global LoRA-FA model and the corresponding expert.
III-B2 MoE-inspired Adaptive Optimization Loss
@@
Current adversarial training algorithms predominantly focus on designing loss functions that balance the relationship between benign data and adversarial samples, seeking an optimal trade-off between accuracy and robustness. We propose a Mixture-of-Experts inspired method for personalized training. This approach adaptively optimizes the model for each client, tailoring the training process to unique data distributions and adversarial characteristics, enhancing adversarial training for each client’s unique local data.
@@
$h_{i_{c}}^{Base}=\frac{1-\gamma}{1-\gamma^{n_{i_{c}}}},$ (10)
@@
$h_{i_{c}}^{Ada}(t)=\epsilon\cdot h_{i_{c}}^{Ada}(t-1)+(1-\epsilon)\cdot h_{i_{c}}^{Base}(t),$ (11)
@@
$\mathcal{L}_{A}(w_{i},x^{adv},y)=\min\sum_{(x,y)\in\mathcal{D}_{i}}h_{i_{y}}\cdot\mathcal{L}_{CE}(\mathcal{F}(w_{i},x^{adv}),y),$ (13)
@@
$\mathcal{L}_{S}(w_{i},x,x^{adv})=\mathcal{L}_{KL}(\mathcal{F}(w_{i},x),\mathcal{F}(w_{i},x^{adv}))$ (14)
@@
The optimization objective for the backbone module of pre-trained models is to enhance its generalization in extracting features from adversarial samples, while the classifier module focuses on achieving personalized classification performance. Inspired by the MoE framework, we propose a collaborative approach, guiding the backbone to shift toward clusters with similar distributions. Each client adjusts its model toward its local cluster, leveraging a fused reference model combining the global model $w_{g}^{L}$ and the expert model $w_{e,i}^{L}$:
$w_{i}^{ref}=\left(1-\eta\right)w_{g}^{L}+\eta w_{e,i}^{L},$ (15)
where $\eta$ is a hyperparameter controlling the fusion ratio. The local training loss is then computed based on the distance between the client’s parameters and the reference model:
$\mathcal{L}_{R}(w_{i}^{L},w_{i}^{ref})=\left\|w_{i}^{L}-w_{i}^{ref}\right\|_{2}^{2},$ (16)
@@
Finally, by combining eq. 13, eq. 14 and eq. 16, we weight the above loss terms to obtain the adversarial training loss as
$\mathcal{L}=\mathcal{L}_{A}+\lambda_{1}\mathcal{L}_{S}+\lambda_{2}\mathcal{L}_{R}.$ (17)
@@
Each client performs local adversarial training using the objective in eq. 17, updating only its LoRA-FA and classifier modules. The LoRA-FA parameters are then periodically uploaded to the server for global aggregation.
III-C Convergence Analysis
To theoretically establish the convergence upper bound of our method, we derive its convergence properties based on our adversarial training. Let $t\in{1,2,\dots,T_{1}}$ denote the index of global aggregation rounds at the server, and introduce an equivalent step size $\zeta>0$, under which we can define as
$g_{i}^{t}\triangleq\frac{w_{g}^{L,t}-w_{i}^{L,t+1}}{\zeta},\quad\hat{g}^{t}\triangleq\sum_{i=1}^{N}\bar{q}_{i}^{t}g_{i}^{t}.$ (18)
The aggregation step can be expressed as $w_{g}^{L,t+1}=w_{g}^{L,t}-\zeta\left.\hat{g}^{t}\right$.. The local optimization can be defined by $\tilde{f}_{i}(w_{i}^{L},w_{i}^{C})$. Accordingly, the personalized LoRA training can be formulated as
$\Phi(w^{L})\triangleq\frac{1}{N}\sum_{i=1}^{N}\min_{w^{C}}\tilde{f}_{i}(w_{i}^{L},w_{i}^{C}).$ (19)
We make the following assumptions:
Assumption 1. $\Phi$ is $\mathrm{L}$-smooth and lower bounded by $\Phi_{\mathrm{inf}}$, defined as
$\|\nabla\Phi(u)-\nabla\Phi(v)\|\leq\mathrm{L}\|u-v\|,\quad\Phi(w)\geq\Phi_{\inf}.$ (20)
Assumption 2. We assume that the normalized weights and update directions contributed by honest clients are unbiased and have bounded variance, which can be expressed as:
$\bar{q}_{i,\mathbf{H}}^{t}\triangleq\frac{\tilde{q}_{i}^{t}|D_{i}|}{\sum_{j\in\mathbf{H}}\tilde{q}_{j}^{t}|D_{j}|},\quad g_{\mathbf{H}}^{t}\triangleq\sum_{i\in\mathbf{H}}\bar{q}_{i,\mathbf{H}}^{t}g_{i}^{t},$ (21)
where $\mathbb{E}[g_{\mathbf{H}}^{t}\mid w_{g}^{L,t}]=\nabla\Phi(w_{g}^{L,t})$, indicating unbiasedness, and $\mathbb{E}{\left[\|g_{\mathbf{H}}^{t}-\nabla\Phi(w_{g}^{L,t})\|^{2}\mid w_{g}^{L,t}\right]}\leq\pi^{2}$.
Assumption 3. The filtered aggregation direction is unbiased with respect to the honest reference, i.e., $\mathbb{E}[\hat{g}^{t}\mid w_{g}^{L,t}]=\mathbb{E}[g_{\mathbf{H}}^{t}\mid w_{g}^{L,t}]$. The second-order residual moment has bounded as $\mathbb{E}\left[\|\hat{g}^{t}-g_{\mathbf{H}}^{t}\|^{2}\mid w_{g}^{L,t}\right]\leq b_{t}^{2}$.
We obtain the following theorem and its proof:
Theorem 1. When $\zeta\leq 1/\mathrm{L}$, we have
$\frac{1}{T_{1}}\sum_{t=0}^{T_{1}-1}\mathbb{E}$ $\|\nabla\Phi(w_{g}^{L,t})\|^{2}\leq\frac{2\left(\Phi(w_{g}^{L,0})-\Phi_{\inf}\right)}{\zeta T_{1}}$ (22)
$+\mathrm{L}\zeta\left(\pi^{2}+\frac{1}{T_{1}}\sum_{t=0}^{T_{1}-1}b_{t}^{2}\right).$
Proof 1. By the descent lemma in Assumption 1, for any $t$, we have
$\Phi(w_{g}^{L,t+1})$ $\leq\langle\nabla\Phi(w_{g}^{L,t}),w_{g}^{L,t+1}-w_{g}^{L,t}\rangle+$ (23)
$\Phi(w_{g}^{L,t})+\frac{\mathrm{L}}{2}\|w_{g}^{L,t+1}-w_{g}^{L,t}\|^{2}.$
By $w_{g}^{L,t+1}=w_{g}^{L,t}-\zeta\hat{g}^{t}$ we get
$\Phi(w_{g}^{L,t+1})\leq\Phi(w_{g}^{L,t})-\zeta\langle\nabla\Phi(w_{g}^{L,t}),\hat{g}^{t}\rangle+\frac{\mathrm{L}\zeta^{2}}{2}\|\hat{g}^{t}\|^{2}.$ (24)
Let $u^{t}=\hat{g}^{t}-\nabla\Phi(w_{g}^{L,t})$, we have $\hat{g}^{t}=\nabla\Phi(w_{g}^{L,t})+u^{t}$ and $\|\hat{g}^{t}\|^{2}=\|\nabla\Phi(w_{g}^{L,t})\|^{2}+2\langle\nabla\Phi(w_{g}^{L,t}),u^{t}\rangle+\|u^{t}\|^{2}$. Under Assumptions 2–3, we get $\mathbb{E}[u^{t}\mid w_{g}^{L,t}]=0$, and thus
$\mathbb{E}[\langle\nabla\Phi(w_{g}^{L,t}),\hat{g}^{t}\rangle\mid w_{g}^{L,t}]=\|\nabla\Phi(w_{g}^{L,t})\|^{2},$ (25)
$\mathbb{E}[\|\hat{g}^{t}\|^{2}\mid w_{g}^{L,t}]=\|\nabla\Phi(w_{g}^{L,t})\|^{2}+\mathbb{E}[\|u^{t}\|^{2}\mid w_{g}^{L,t}].$ (26)
Taking expectation on the above expression and substituting it back into the preceding descent inequality yields
$\mathbb{E}[\Phi(w_{g}^{L,t+1})]\leq\mathbb{E}[\Phi(w_{g}^{L,t})]$- $\left(\zeta-\frac{\mathrm{L}\zeta^{2}}{2}\right)\mathbb{E}\|\nabla\Phi(w_{g}^{L,t})\|^{2}$ (27)
$+\frac{\mathrm{L}\zeta^{2}}{2}\mathbb{E}\|u^{t}\|^{2}$
Furthermore, since $u^{t}=(\hat{g}^{t}-g_{\mathbf{H}}^{t})+(g_{\mathbf{H}}^{t}-\nabla\Phi(w_{g}^{L,t}))$, and using the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$, Assumptions 2-3 imply $\mathbb{E}\|u^{t}\|^{2}\leq 2b_{t}^{2}+2\pi^{2}$. Finally, choosing $\zeta\leq 1/\mathrm{L}$ yeilds $\zeta-\frac{\mathrm{L}\zeta^{2}}{2}\geq\frac{\zeta}{2}$. Summing over $t=0,\ldots,T_{1}-1$ and applying a telescoping sum together with $\Phi(w)\geq\Phi_{\inf}$ completes the proof of the theorem.
III-D Forward-gating Selection Fine-tuning